A biker travels 400m E then turns 300m S before going 200m W then 250m NW. What is the resultant vector using the component method?


1
Expert's answer
2021-10-01T13:27:34-0400

Find the East and North component of each displacement:


"A_E=400\text{ m},\\
A_N=0\text{ m}.\\
B_E=0\text{ m},\\
B_N=-300\text{ m}.\\
C_E=-200\text{ m},\\
C_N=0\text{ m}.\\
D_E=-250\sin45°=-177\text{ m},\\
D_N=250\sin45°=177\text{ m}.\\\space\\
R_E=A_E+B_E+C_E+D_E=23\text{ m},\\
R_N=A_N+B_N+C_N+D_N=-122\text{ m},\\
R=\sqrt{R_N^2+R_E^2}=124\text{ m},\\\space\\
\theta=\arctan\frac{R_N}{R_E}=-79.3°\text{ from E to N}"